Investments Millennium Securities last year invested $600000 in mutual funds. The International fund earned 11% Simple interest. The high-tech fund earned 7% simple interest. At the end of the year, the company had earned $50000 in interest. how much did it invest in each fund?
Answer by checkley77(12844)

.11x+.07(600,000-x)=50,000
.11x+42,000-.07x=50,000
.04x=50,000-42,000
.04x=8,000
x=8,000/.04
x=$200,000 amount invested @ 11%.
600,000-200,000=$400,000 amount invested @ 7%.
Proof:
.11*200,000+.07*400,000=50,000
22,000+28,000=50,000
50,000=50,000